if angle x is acute and sin x=7/25 find cos(180-x)
Answer by Alan3354(69443)

if angle x is acute and sin x=7/25 find cos(180-x)
------------
cos(x) = sqrt(1 - sin^2(x)) = 24/25
cos(180 - x) --> Q2
cos(180 - x) = -24/25